# arXiv:2603.01498v1[cs.CV] 02 Mar 2026

### v1

@@ 4.5 Ablation Study
We conducted systematic ablation experiments on the Gaza-change and SECOND datasets to validate the effectiveness of the proposed third path and the MLHA module. The experimental results are shown in Table 3 and Table 4, where Path indicates whether the third path is introduced, and MLHA indicates whether the Multi-Level Hybrid Attention module is used in the decoder. Evidently, the performance when using either the MLHA module alone or introducing the third path (Path) alone is significantly better than the baseline, where neither module is used. This demonstrates the important contribution of both multi-level attention fusion and fine-grained complementary feature modeling to improving MCD performance. When both Path and MLHA are enabled (i.e., the complete Tri-path DINO), the model achieves the best performance across all key metrics on both datasets. This indicates a significant synergistic and complementary relationship between the encoder and decoder module designs.
Table 3: Module-level ablation results on the Gaza-change dataset. Best scores are in bold.
@@
Table 4: Module-level ablation results on the SECOND dataset. Best scores are in bold.

### v2

@@ IV-E Ablation Study
We conducted systematic ablation experiments on the Gaza-change and SECOND datasets to validate the effectiveness of the proposed third path and the MLHA module. The experimental results are shown in Table III and Table IV, where Path indicates whether the third path is introduced, and MLHA indicates whether the Multi-Level Hybrid Attention module is used in the decoder. Evidently, the performance when using either the MLHA module alone or introducing the third path (Path) alone is significantly better than the baseline, where neither module is used. This demonstrates the important contribution of both multi-level attention fusion and fine-grained complementary feature modeling to improving MCD performance. When both Path and MLHA are enabled (i.e., the complete Tri-path DINO), the model achieves the best performance across all key metrics on both datasets. This indicates a significant synergistic and complementary relationship between the encoder and decoder module designs.
TABLE III: Module-level ablation results on the Gaza-change dataset. Best scores are in bold.
@@
TABLE IV: Module-level ablation results on the SECOND dataset. Best scores are in bold.